# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v4

@@
New rotating dilaton black hole and black string solutions in three spacetime dimensions are obtained. The background spacetime interpolates between Anti-de Sitter and a (an asymptotically) flat spacetime. The new black strings are characterized by their masses, angular momenta and axion charges. The uncharged black string solutions have only a single horizon. They contain a curvature singularity enclosed inside their horizons. Thus, they are also black hole solutions of Einstein-scalar gravity. On the other hand, the charged black string solutions have two horizons. They may or may not contain a curvature singularity depending on the ratio of their inner and outer horizons radii. When they contain a singularity, the singularity is either at or enclosed inside their inner horizons. We also obtain, in addition to the uncharged black strings, other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black strings. The new black holes are described by their masses and angular momenta. We show that their angular momenta do not depend on their horizons radii. We also find that the masses of a class of the black holes can be made negative by adjusting, using the gauge ambiguity, the asymptotic value of the (independent component of the) Kalb-Ramond field in the dual black strings. We also discuss black hole and black string solutions with a curvature singularity at or beyond their outer or event horizons. Black hole solutions with a ring curvature singularity in between their inner and outer horizons are also presented.
@@ 1. Introduction
In this paper, we obtain new black hole and black string solutions in the string background obtained in [1]. The string background has the product topology ${\cal A}_{3}\times S^{3}\times{\cal X}_{4}$ where $S^{3}$ denotes a unit three-sphere and ${\cal X}_{4}$ is a compact four dimensional manifold. Examples include the Calabi-Yau manifolds $T^{4}$ and $K3$. The spacetime ${\cal A}_{3}$ interpolates between $AdS_{3}$ in the infrared and a (an asymptotically) linear dilaton spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$ in the ultraviolet. The worldsheet conformal field theory on ${\cal A}_{3}$ is an exact marginal current bilinear deformation of the worldsheet theory on $AdS_{3}$. In the boundary theory the deformation is equivalent to a deformation by an irrelevant operator of (left and right) dimensions $(2,2)$, see [[1]1,,  [2]2] and references therein. The resulting deformed theory is thus non-conformal and dual to string theory on ${\cal A}_{3}$ (times the internal product manifold ${\cal M}_{7}:=S^{3}\times{\cal X}_{4}$).
In the boundary theory the deformation coupling $\widehat{\gamma}$ has (left and right) mass dimensions $(-1,-1)$. The coupling can have either signs. In the case it is negative the deformation washes the holographic boundary away to infinity. On the other hand, in the case it is positive, in general it moves the boundary into the bulk. In Anti-de Sitter (AdS) spacetime moving the boundary into the bulk generates at onset a curvature singularity. For more and thorough discussions see [1]. Therefore, we will make in our following discussions a distinction between positive and negative coupling. Also since products of conformal field theories (CFTs) are still CFTs, we will simply ignore the (worldsheet) CFT on the internal manifold ${\cal M}_{7}$.
@@
In three spacetime dimensions the Hodge dual of the Kalb-Ramond field is a (pseudo-)scalar which is identified as an axion field [6]. Its value at infinity gives an additional parameter to describe the solutions.2 T-duality relates it to angular momentum. The solutions we find are asymptotically flat (in the string frame) and they are exact solutions of classical string theory; we will be precise later in the paper.
@@
The new black string solutions are parametrized by their masses, angular momenta and axion charges. The axion charge is the value of the (pseudo-)scalar at the boundary or spatial infinity. In general it is determined by the electric flux in the world-volume of the black string. We show that the axion charge is given only in terms of the deformation coupling. In particular it does not depend on the horizons (radii). The uncharged solutions, i.e., those black string solutions with zero axion charge, are also black hole solutions of Einstein-scalar gravity. We also obtain other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black string solutions.3 See Appendix B for a review of the duality transformation. We show that their angular momenta are given by the dual black strings axion charges and thus do not depend on their horizons. We also show that the masses of a class of the black hole solutions can be made negative by changing (in the dual black strings) the value of the (component of the) Kalb-Ramond field at radial infinity.
@@
The uncharged black strings have only a single horizon. They contain a curvature singularity with the structure of a ring. In the case the deformation coupling is negative the singularity is enclosed inside the event horizon.4 This particular class of solutions, i.e., the class of solutions characterized by zero axion charge and negative coupling (of $-1$), is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a scalar field. However, here we simply obtain it using duality and coordinate transformations from global $AdS_{3}$. We will discuss it in more detail later in the paper. Otherwise, i.e., for positive coupling, the singularity is beyond the event horizon. The charged black strings, on the other hand, possess two horizons. We show that depending on the ratio of the two horizons radii the solutions may or may not contain a curvature singularity. The singularity has the structure of a ring. In the case the deformation coupling is negative the singularity is at or enclosed inside the inner horizon.5 The inner (event) horizon is also referred to as Cauchy horizon. Otherwise, it is at or beyond the outer horizon. The new black hole solutions that we obtain by applying duality transformation have only a single horizon. They contain a curvature singularity. The singularity is enclosed inside the event horizon. To end, the black hole solutions that we will discuss, including those characterized by positive coupling, have only a single horizon and a ring singularity. The event horizon divides the spacetime into two regions. The interior region contains the singularity and the exterior region contains the rest of the spacetime. This will be our definition of interior and exterior. In the case the coupling is positive the ring singularity cannot be contracted to the (coordinate) origin without crossing the horizon(s).
@@
The paper is organized as follows. In section two we review the rotating BTZ black hole [[8]8,,  [9]9]. In section three we obtain the new black hole and black string solutions. In section four we discuss the dilaton black hole solutions. These are black hole solutions of the Einstein-scalar gravity. We compute their ADM masses and angular momenta. In section five we summarize the main results and discuss future research directions. In Appendix A we review a parameterization of $AdS_{3}$ that in particular is better suited in relation to BTZ black hole. In Appendix B we review the target-space duality, also known as T-duality, transformation which we use in section three. In Appendix C we review the Brown-York quasi-local energy and mass. In Appendix D we review Kruskal approach which we use to remove coordinate singularities and obtain Penrose diagrams for the solutions we consider in the paper.
@@ 2. BTZ black hole
where $t,\omega\in{\tenrm I\kern-1.79993ptR}$, and extend the domain of $r$ to all the real numbers, i.e., $r={\overline{r}}\in{\tenrm I\kern-1.79993ptR}$, see A.5. We define, since we are in particular interested in the region exterior to the black hole horizons, $\rho$ as
@@
where $\rho\geq\rho_{+}>\rho_{-}$. In our convention the coordinates $\rho$ and $t$ have mass dimension $-1$. Using the $GL(2)$ transformation (2.4) and the change of variable (2.5) in (2.3) we get the metric
@@
The metric is invariant under $\rho\to-\rho$. The BTZ black hole metric is obtained after identifying $\omega$ periodically. In the metric (2.6) we now identify $\omega$ with period $2\pi$, i.e., $\omega\sim\omega+2\pi$. This in turn implies $(\overline{\varphi},\overline{\psi})\sim(\overline{\varphi}+2\pi\rho_{+}/l,\overline{\psi}+2\pi\rho_{-}/l)$. Now $\rho$ is radial coordinate, i.e. $\rho\geq 0$. Therefore, BTZ black hole can be viewed as a quotient space of $AdS_{3}$. Thus, locally it is isometric to $AdS_{3}$. Furthermore, the scalar curvature is constant. It is given by8 In $AdS_{d}$ the scalar curvature and cosmological constant are given by $R=-{d(d-1)\over l^{2}},\quad\Lambda=-{(d-1)(d-2)\over 2l^{2}}.$ (2.7)2.7( 2.7 )
@@
where $\Lambda$ is the cosmological constant. It approaches (without identifications) $AdS_{3}$ spacetime asymptotically.
@@
in which the Killing vector field $(\partial_{t})^{\mu}=(1,0,0):=\delta^{\mu}_{t}$ is space-like defines an ergosphere.9 The analytically extended BTZ metric is obtained by replacing $\rho$ with $i\rho$ or simply by setting $\rho^{2}=r\in{\tenrm I\kern-1.79993ptR}$. It contains closed time-like curves and thus usually it is not considered.
The BTZ black hole metric (2.6) satisfies the vacuum Einstein equations in $2+1$ spacetime dimensions,
@@
Thus, it is described by two parameters, its mass $M$ and angular momentum $J$. We will obtain $M$ and $J$ later in the paper following the Brown-York approach. In order for the solution to describe black hole or for a horizon to exist we assume
@@
We note that $(M,J)=(-1,0)$ is the $AdS_{3}$ metric (2.3). It is separated from the vacuum solution, i.e., $(M,J)=(0,0)$, or the continuous black hole spectrum, by a mass unit. The solutions with $-1<M<0,J=0$ have naked conical singularity at the origin $\rho=0$. For small $\rho$ the metric is
@@
Thus, it describes point particle sources with negative cosmological constant [[15]15,,  [16]16].11 See [17] for numerical solutions of static non-singular vortices in $AdS_{3}$ that are not BTZ black holes. The zero mass black hole is identified in $(1,1)$ $AdS$ supergravity [18] as the Ramond sector ground state. The Neveu-Schwarz sector ground state is global $AdS_{3}$.
In the next section we obtain a family of rotating black hole and black string solutions by applying a similar $GL(2)$ transformation since in the infrared the spacetime ${\cal A}_{3}$ is well described by $AdS_{3}$. We explain this in the next section.
@@ 3.1. New black string solutions
where $g_{s}$ is the string coupling at $\gamma=0$. Note the absolute value sign is necessary in the case $\gamma$ is positive to ensure the dilaton is real everywhere. The Kalb-Ramond field strength $H$ is given by
@@
The time coordinate $\varphi$ takes its value in ${\tenrm I\kern-1.79993ptR}$. The radial coordinates $\theta$ takes its value in ${\tenrm I\kern-1.79993ptR}^{+}$, i.e. $0\leq\theta<\infty$. The angular variable $\psi$ is periodic. It has period $2\pi(1-\gamma)$ to ensure that there is no conical singularity. The deformation coupling $\gamma$ takes the values in the range
@@
$AdS_{3}$ (with Kalb-Ramond two form) is an exact solution to (classical) string theory, i.e., there are no $\alpha^{\prime}=l_{s}^{2}$ corrections to $AdS_{3}$. $l_{s}$ is the string length. The solution ${\cal A}_{3}$ is related to $AdS_{3}$ by T-dualities and coordinate shift (see [1] for details), and thus it is exact in $\alpha^{\prime}$. However, it is perturbative in $g_{s}$ [19]. On the world-sheet the deformation corresponds to an exact marginal deformation (see, e.g., [1]). More on this in a future work. Later in the section we also give previously known $\alpha^{\prime}$ exact (black string) solutions (in ${\cal A}_{3}$) as special cases.
@@
The time coordinate $t$ takes its value in ${\tenrm I\kern-1.79993ptR}$. We will in general assume $x$ is an angle coordinate. $\rho$ is a radial coordinate. We will in general assume, unless stated otherwise, that $\rho_{-}<\rho_{+}$. The periodicity of $x$ is chosen so that there is no conical singularity at the origin. We note that at $\rho=\infty$ the $B$ field, i.e., $B_{tx}$, is constant. We can adjust the constant using the gauge freedom. This has, as we will show in the next section, a non-trivial consequence. In particular, the constant enters into physical observables.

### v5

@@
New rotating dilaton black hole and black string solutions in three spacetime dimensions are obtained. The background spacetime interpolates between Anti-de Sitter and a (an asymptotically) flat spacetime. The new black strings are characterized by their masses, angular momenta and axion charges. The uncharged black string solutions have only a single horizon. Enclosed inside their horizons, they contain a curvature singularity. Thus, they are also black hole solutions of Einstein-scalar gravity. On the other hand, the charged black string solutions have two horizons. They may or may not contain a curvature singularity depending on the ratio of their inner and outer horizons radii. When they contain a singularity, the singularity is either at or enclosed inside their inner horizons. We also obtain, in addition to the uncharged black strings, other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black strings. The new black holes are described by their masses and angular momenta. We show that their angular momenta do not depend on their horizons radii. We also find that the masses of a class of the black holes can be made negative by adjusting, using the gauge ambiguity, the asymptotic value of the (independent component of the) Kalb-Ramond field in the dual black strings. We also discuss black hole and black string solutions with a curvature singularity at or beyond their outer or event horizons. Novel black hole solutions with a ring curvature singularity in between their inner and outer horizons are also presented.
@@ 1. Introduction
In this paper, we obtain new black hole and black string solutions in the string background obtained in [1]. The string background has the product topology ${\cal A}_{3}\times S^{3}\times{\cal X}_{4}$, where $S^{3}$ denotes a unit three-sphere and ${\cal X}_{4}$ is a compact four dimensional manifold. Examples of ${\cal X}_{4}$ include the Calabi-Yau manifolds $T^{4}$ and $K3$. The spacetime ${\cal A}_{3}$ interpolates between $AdS_{3}$ in the infrared and a (an asymptotically) linear dilaton spacetime ${\tenrm I\kern-1.79993ptR}\times S^{1}\times{\tenrm I\kern-1.79993ptR}$ in the ultraviolet. The world-sheet conformal field theory on ${\cal A}_{3}$ is an exact marginal current bilinear deformation of the world-sheet theory on $AdS_{3}$. In the boundary theory the deformation is equivalent to a deformation by an irrelevant operator of (left and right) scaling dimensions $(2,2)$, see [[1]1,,  [2]2] and references therein. The resulting deformed theory is thus non-conformal and dual to string theory on ${\cal A}_{3}$ (times the internal product manifold ${\cal M}_{7}:=S^{3}\times{\cal X}_{4}$).
In the boundary theory the deformation coupling $\widehat{\gamma}$ has mass dimension $-2$. The coupling can have either signs. In the case it is negative the deformation washes the holographic boundary away to infinity. On the other hand, in the case it is positive, in general it moves the boundary into the bulk. In Anti-de Sitter ($AdS$) spacetime moving the boundary into the bulk generates at onset a curvature singularity. For more and thorough discussions see [1]. Therefore, we will make in our following discussions a distinction between positive and negative coupling. Also since products of conformal field theories (CFTs) are still CFTs, we will simply ignore the (world-sheet) CFT on the internal manifold ${\cal M}_{7}$. Therefore, ${\cal M}_{7}$ is unaffected throughout our discussion.
@@
In three spacetime dimensions the Hodge dual of the Kalb-Ramond field is a (pseudo-)scalar which is identified as an axion field [6]. Its value at infinity gives an additional parameter to describe the solutions.2 T-duality relates it to angular momentum. The solutions we find are asymptotically flat (in the string frame) and they are exact solutions of classical string theory.
@@
The new black string solutions are parametrized by their masses, angular momenta and axion charges. The axion charge is the value of the (pseudo-)scalar at the boundary or spatial infinity. In general it is determined by the electric flux in the world-volume of the black string. We show that the axion charge is given only in terms of the deformation coupling. In particular it does not depend on the horizons (radii). The uncharged solutions, i.e., those black string solutions with zero axion charge, are also black hole solutions of Einstein-scalar gravity. We also obtain other new black hole solutions of Einstein-scalar gravity by applying duality transformation on the new charged black string solutions.3 See Appendix B for a review of the duality transformation. We show that their angular momenta are given by the dual black strings axion charges and thus do not depend on the horizons. We also show that the masses of a class of the black hole solutions can be made negative by changing (in the dual black strings) the value of the (component of the) Kalb-Ramond field at radial infinity.
@@
The uncharged black strings have only a single horizon. They contain a curvature singularity with the structure of a ring. In the case the deformation coupling is negative the singularity is enclosed inside the event horizon.4 This particular class of solutions, i.e., the class of solutions characterized by zero axion charge and negative coupling (of $-1$), is already known in the literature [7]. It is obtained by directly solving the Einstein field equations with a minimally coupled scalar field. However, here we simply obtain it using duality and coordinate transformations from global $AdS_{3}$. We will discuss it in more detail later in the paper. Otherwise, i.e., for positive coupling, the singularity is beyond the event horizon. The charged black strings, on the other hand, possess two horizons. We show that depending on the ratio of the two horizons radii the solutions may or may not contain a curvature singularity. The singularity has the structure of a ring. In the case the deformation coupling is negative the singularity is at or enclosed inside the inner horizon.5 The inner (event) horizon is also referred to as Cauchy horizon. Otherwise, it is at or beyond the outer horizon. The new black hole solutions that we obtain by applying duality transformation have only a single horizon. They contain a curvature singularity. The singularity is enclosed inside the event horizon. To end, the black hole solutions that we will discuss, including those characterized by positive coupling, have only a single horizon and a ring singularity. The event horizon divides the spacetime into two regions. The interior region contains the singularity and the exterior region contains the rest of the spacetime. This will be our definition of interior and exterior. In the case the coupling is positive the ring singularity cannot be contracted to the (coordinate) origin without crossing the horizon(s).
@@
The paper is organized as follows. In section two we review the rotating BTZ black hole [[8]8,,  [9]9]. In section three we obtain the new (dilaton) black hole and black string solutions. In section four we further discuss the new (dilaton) black hole solutions. They are (black hole) solutions of a minimally coupled Einstein-scalar gravity theory. We in particular compute their ADM masses and angular momenta. In section five we summarize the main results and discuss future research directions. We also give, in section five, novel black hole solutions with a ring curvature singularity in between their inner and outer horizons.
In Appendix A we review a parameterization of $AdS_{3}$ that in particular is better suited in relation to BTZ black hole. In Appendix B we review T-duality, also known as target-space duality. We use the T-duality transformation in section three. In Appendix C we review the Brown-York quasi-local energy and mass. In Appendix D we review the Kruskal approach. We use it to remove the coordinate singularities of the solutions we study in the paper and to obtain their Penrose diagrams.
@@ 2. BTZ black hole
where $t,\omega\in{\tenrm I\kern-1.79993ptR}$, and extend the domain of $r$ to all the real numbers, i.e., $r={\overline{r}}\in{\tenrm I\kern-1.79993ptR}$, see A.5. We define, since we are in particular interested in the region exterior to the black hole horizons, the new coordinate $\rho$ as
@@
where $\rho\geq\rho_{+}>\rho_{-}$. In our convention the coordinates $\rho$ and $t$ have mass dimension $-1$. Using both the $GL(2)$ transformation (2.4) and the change of variable (2.5) in (2.3) we get the metric
@@
The metric is invariant under $\rho\to-\rho$. The BTZ black hole metric is obtained after identifying $\omega$ periodically. In the metric (2.6) we now identify $\omega$ with period $2\pi$, i.e., $\omega\sim\omega+2\pi$. This in turn implies $(\overline{\varphi},\overline{\psi})\sim(\overline{\varphi}+2\pi\rho_{+}/l,\overline{\psi}+2\pi\rho_{-}/l)$. Now $\rho$ is a radial coordinate, i.e. $\rho\geq 0$. Therefore, BTZ black hole can be viewed as a quotient space of $AdS_{3}$. Thus, locally it is isometric to $AdS_{3}$. Furthermore, the scalar curvature is constant. It is given by8 In $AdS_{d}$ the scalar curvature and cosmological constant are given by $R=-{d(d-1)\over l^{2}},\quad\Lambda=-{(d-1)(d-2)\over 2l^{2}}.$ (2.7)2.7( 2.7 )
@@
where $\Lambda$ is the cosmological constant. Asymptotically, the BTZ black hole approaches (without identifications) $AdS_{3}$ spacetime.
@@
in which the Killing vector field $(\partial_{t})^{\mu}=(1,0,0):=\delta^{\mu}_{t}$ is space-like defines an ergosphere.9 The analytically extended BTZ metric is obtained by replacing $\rho$ with $i\rho$ or simply by setting $\rho^{2}=r\in{\tenrm I\kern-1.79993ptR}$. It contains closed time-like curves and thus usually it is not considered. $r=0$ is a causal singularity.
The BTZ black hole metric (2.6) satisfies the vacuum Einstein equations in $(2+1)$ spacetime dimensions,
@@
Thus, it is described by two parameters, its mass $M$ and angular momentum $J$. We will obtain $M$ and $J$ later in the paper following the Brown-York approach. In order for the solution to describe a black hole or for a horizon to exist we assume
@@
We note that $(M,J)=(-1,0)$ is the $AdS_{3}$ metric (2.3). It is separated from the massless BTZ solution, i.e., $(M,J)=(0,0)$, or the continuous black hole spectrum, by a mass unit. The solutions with $-1<M<0,J=0$ have naked conical singularity at the origin $\rho=0$. For small $\rho$ the metric is
@@
Thus, it describes point particle sources with negative cosmological constant [[15]15,,  [16]16].11 See [17] for numerical solutions of static non-singular vortices in $AdS_{3}$ that are not BTZ black holes. The zero mass black hole is identified in $(1,1)$ $AdS$ supergravity [18] as the Ramond sector ground state. Global $AdS_{3}$ is identified as the Neveu-Schwarz sector ground state.
In the next section we obtain a new family of rotating black hole and black string solutions by applying to ${\cal A}_{3}$, since in the infrared the spacetime ${\cal A}_{3}$ is well described by $AdS_{3}$, the same $GL(2)$ transformation. We explain this in the next section.
@@ 3.1. New black string solutions
where $g_{s}$ is the string coupling at $\gamma=0$. Note the absolute value sign is necessary in the case $\gamma$ is positive to ensure the dilaton is real everywhere. More on this momentarily. The Kalb-Ramond field strength $H$ is given by
@@
The time coordinate $\varphi$ takes its value in ${\tenrm I\kern-1.79993ptR}$. The radial coordinates $\theta$ takes its value (in the case $\gamma\leq 0$) in ${\tenrm I\kern-1.79993ptR}^{+}$, i.e. $0\leq\theta<\infty$. The angular variable $\psi$ is periodic. It has period $2\pi(1-\gamma)$ to ensure that there is no conical singularity at $\theta=0$. The deformation coupling $\gamma$ takes the values in the range
@@
In a moment, we replace $\theta$ using a change of variable by a function of a new coordinate $\rho$. When $\gamma>0$, the range of the radial coordinate $\rho$ is restricted so that the argument of the absolute value (3.4) is always positive. Therefore, in this paper one may simply ignore the absolute value sign. For a detailed discussion on the range of $\theta$ when $\gamma>0$, see [1].
$AdS_{3}$ (with Kalb-Ramond two form) is an exact solution to (classical) string theory, i.e., there are no $\alpha^{\prime}=l_{s}^{2}$ corrections to $AdS_{3}$. $l_{s}$ is the string length. The solution ${\cal A}_{3}$ is related to $AdS_{3}$ by (abelian) T-dualities and coordinate shift (see [1] for details), and thus it is exact in $\alpha^{\prime}$. However, it is perturbative in $g_{s}$ [19]. In general, an abelian T-duality maps an exact solution of string theory into an exact one (see, e.g., [20]). However, in the case the solution is an effective one, abelian T-duality can generate $\alpha^{\prime}$ corrections [21]. Later in the section we give previously known $\alpha^{\prime}$ exact (black string) solutions as special cases of our results.
On the world-sheet the deformation corresponds to an exact marginal deformation (see, e.g., [1]). More on this in a future work.
@@
We will shortly explain why we refer to the solution as black string. The time coordinate $t$ takes its value in ${\tenrm I\kern-1.79993ptR}$. We will in general assume $x$ is an angle coordinate. $\rho$ is a radial coordinate. We will in general also assume, unless stated otherwise, that $\rho_{-}<\rho_{+}$. The periodicity of $x$ is chosen so that there is no conical singularity at the origin. We note that at $\rho=\infty$ the $B$ field, i.e., $B_{tx}$, is constant. We can adjust the constant using the gauge freedom. This has, as we will show in the next section, a non-trivial consequence. In particular, the constant enters into physical observables.